OpenStudy (anonymous):

Solve each equation. Check your answers. 23. r-17=7 24. 180=220-k 25. 74-l=59

OpenStudy (gw2011):

Solve: 23) r=7+17=24; r=24 24) k=220-180=40; k=40 25) i=74-59=15; i=15 Check: 23) 24-17=7 24) 180=220-40 25) 74-15=59